# Provably Finding a Hidden Dense Submatrix among Many Planted Dense Submatrices via Convex Programming

### v1

@@ Abstract
We consider the densest submatrix problem, which seeks the submatrix of fixed size of a given binary matrix that contains the most nonzero entries. This problem is a natural generalization of fundamental problems in combinatorial optimization, e.g., the densest subgraph, maximum clique, and maximum edge biclique problems, and has wide application the study of complex networks. Much recent research has focused on the development of sufficient conditions for exact solution of the densest submatrix problem via convex relaxation. The vast majority of these sufficient conditions establish identification of the densest submatrix within a graph containing exactly one large dense submatrix hidden by noise. The assumptions of these underlying models are not observed in real-world networks, where the data may correspond to a matrix containing many dense submatrices of varying sizes.
@@ 1 Introduction
We consider the densest matrix problem (DSM), in which one aims to identify the submatrix of given size containing a maximum number of nonzero entries; see Lanciano et al. (2024) for a survey of the historical and recent study of the DSM. The densest submatrix can be considered a generalization of the densest bipartite subgraph problem: given a bipartite graph $G=(V,U,E)$ and integers $k_{1}$ and $k_{2}$, the objective is to extract a subgraph induced by $k_{1}$ nodes from $V$ and $k_{2}$ nodes from $U$ that maximizes the number of edges. This further generalizes the densest subgraph problem: given graph $G$ and integer $k$, find the $k$-node subgraph containing maximum number of edges.
@@ 1.1 Contributions
In this work, the phrase “with high probability” (abbreviated as w.h.p.) refers to events whose probability converges to 1 as the problem size increases. Formally, for a sequence of random events $\{E_{n}\}$, we say that $E_{n}$ occurs with high probability if
$\Pr(E_{n})\to 1\quad\text{as }n\to\infty.$
polynomially as $n\to\infty$. Equivalently, $\Pr(E_{n})\geq 1-1/p(n)$ for some polynomial function $p(n)$ such that for any $\varepsilon>0$, there exists an integer $N_{\varepsilon}$ such that for all $n\geq N_{\varepsilon}$, we have $p(n)<\epsilon 0$.
@@
for each $i\in U_{r}$, $j\in V_{s}$, $r\in\{1,2,\dots,k_{1}\}$ and $r,s\in\{1,2,\dots,k_{2}\}$.
@@
$\tilde{\sigma}_{*}^{2}:=\frac{p^{*}}{1-p^{*}}\text{ where }p^{*}:=\max_{(r,s)\neq(1,1)}\frac{p_{rs}}{1-p_{rs}}.$
@@
$p_{11}-p^{*}$ $\geq c\cdot\max\left\{\sqrt{\frac{\tilde{\sigma}_{*}^{2}N\log N}{m^{2}}},\sqrt{\frac{\sigma_{11}^{2}\log N}{m}},\frac{\log N}{m}\right\}.$ (1b)
@@
This result generalizes to the unbalanced planted submatrix model, as well as adversarially constructed matrices. We will motivate the relevant convex relaxation and provide these sufficient conditions in following section.
@@ 2.1 Formulation and Relaxation
We define $M]:=\{1,2,...,M\}$ and $N]:=\{1,2,...,N\}$ for every positive integer $M$ and $N$ respectively. Let $\bar{U}_{1}\subseteq[M]$ and $\bar{V}_{1}\subseteq[N]$ denote the true row and column indices of the planted submatrix. Furthermore, we define vectors $\bar{\boldsymbol{u}}\in\{0,1\}^{M}$ and $\bar{\boldsymbol{v}}\in\{0,1\}^{N}$ as the indicator vectors of the planted row and column sets. The planted submatrix support is given by matrix $\bar{\boldsymbol{X}}=\bar{\boldsymbol{u}}\bar{\boldsymbol{v}}^{\top}\in\{0,1\}^{M\times N}$, e. By construction, $\bar{\boldsymbol{X}}=\bar{\boldsymbol{u}}\bar{\boldsymbol{v}}^{\top}$ is a rank-one binary matrix supported on $\bar{U}_{1}\times\bar{V}_{1}$), such that $\bar{X}_{ij}=1$ if $(i,j)\in\bar{U}_{1}\times\bar{V}_{1}$ and $\bar{X}_{ij}=0$ otherwise.
Suppose that $\Omega$ encodes the set of zero entries in a given matrix $\boldsymbol{A}\in\mathbb{R}^{M\times N}$. Without loss of generality, we may assume that $\boldsymbol{A}$’s entries are binary. If not, we can substitute $\boldsymbol{A}$ with a binary matrix that features the same sparsity pattern, preserving the index set of the densest $m_{1}\times n_{1}$-submatrix. Our objective is to derive a rank-one matrix $\boldsymbol{X}$ with $m_{1}n_{1}$ nonzero entries that minimizes the number of disagreements with $\boldsymbol{A}$ on $\Omega$.
@@
$d(\bar{\boldsymbol{X}})=\frac{1}{m_{1}n_{1}}\left(m_{1}n_{1}-\|P_{\Omega}(\bar{\boldsymbol{X}})\|_{0}\right),$
@@
$\begin{array}[]{cl}\displaystyle\min_{\boldsymbol{X},\boldsymbol{Y}\in\{0,1\}^{M\times N}}&\mathrm{Tr}(\boldsymbol{Y}\boldsymbol{1}\boldsymbol{1}^{T})\\ \mathrm{s.t.}&\mathrm{Tr}(\boldsymbol{X}\boldsymbol{1}\boldsymbol{1}^{T})=m_{1}n_{1},~P_{\Omega}(\boldsymbol{X}-\boldsymbol{Y})=\boldsymbol{0},~\mathrm{rank}(\boldsymbol{X})=1.\end{array}$ (2)
@@
Relaxing both binary and rank constraints gives the convex relaxation:
$\begin{array}[]{cl}\min&\lVert\boldsymbol{X}\rVert_{*}+\gamma\mathrm{Tr}(\boldsymbol{Y}\boldsymbol{1}\boldsymbol{1}^{T})\\ \mathrm{s.t.}&\mathrm{Tr}(\boldsymbol{X}\boldsymbol{1}\boldsymbol{1}^{T})=m_{1}n_{1},~P_{\Omega}(\boldsymbol{X}-\boldsymbol{Y})=\boldsymbol{0},~\boldsymbol{0}\leq\boldsymbol{X}\leq\boldsymbol{1}\boldsymbol{1}^{T},~\boldsymbol{0}\leq\boldsymbol{Y}.\end{array}$ (3)
In the expression, $\gamma\geq 0$ denotes a regularization parameter, selected to balance between the two objectives. As previously mentioned, it is reasonable to anticipate retrieving the solution of (2) from (3) when matrix $\boldsymbol{A}$ consists of a dense $m_{1}\times n_{1}$ submatrix hidden among perhaps many other dense submatrices under certain assumptions about the size and density of these submatrices. In particular, this program succeeds in recovering the planted block exactly when embedded in a noisy binary matrix under the conditions of Theorem 1.1.
@@ 2.2 A Sufficient Condition for Perfect Recovery for Random Matrices
Let $\boldsymbol{A}\in\mathbb{R}^{M\times N}$ be sampled from the planted submatrix model with row partitions $(U_{1},\ldots,U_{k_{1}})$ and column partitions $(V_{1},\ldots,V_{k_{2}})$. Let $m_{i}=|U_{i}|$ and $n_{i}=|V_{i}|$, and let $p_{rs}$ denote the probability that $A_{ij}=1$ for $i\in U_{r}$, $j\in V_{s}$.
@@
$\tilde{\sigma}_{*}^{2}:=\max_{(r,s)\neq(1,1)}\left\{\min\left\{\frac{1-p_{rs}}{p_{rs}},\frac{p_{rs}}{1-p_{rs}}\right\}\right\}.$
Then there exist constants $c_{1},c_{2}>0$ such that the planted block $(U_{1},V_{1})$ is the unique densest $m_{1}\times n_{1}$ submatrix of $\boldsymbol{A}$ and $\boldsymbol{X}^{*}=\boldsymbol{u}_{1}\boldsymbol{v}_{1}^{T}$ unique optimal solution of the convex relaxation (3) if the following conditions hold for $(r,s)\neq(1,1)$
@@
$p_{11}-p_{rs}$ $\geq c_{1}\cdot\max\left\{\sqrt{\frac{\tilde{\sigma}_{*}^{2}N^{*}\log N^{*}}{m_{1}n_{1}}},\sqrt{\frac{\sigma_{11}^{2}\log N^{*}}{n_{1*}}},\frac{\log N^{*}}{n_{1*}}\right\}$ (4c)
@@
$\gamma\in\left(\frac{2}{(p_{11}-p^{*})\sqrt{m_{1}n_{1}}},\frac{c_{2}}{(p_{11}-p^{*})\sqrt{m_{1}n_{1}}}\right),$
@@ 2.3 A Sufficient Condition for Perfect Recovery for Adversarially Generated Matrices
The sufficient conditions for perfect recovery given in Section 2.2 may be overly conservative. For example, the assumption that the planted submatrix of interest is smaller than all other blocks in the generative model is kept largely to ensure that this is the densest $m_{1}\times n_{1}$-submatrix with high probability if the hypothesis of Theorem 2.1 is satisfied. This simplifies the statement of the sufficient conditions, as well as their proof, but is not necessary, in general, to ensure perfect recovery. In this section, we provide a sufficient condition for recovery that holds for deterministically generated graphs. This condition deterministically encodes the relationship between planted submatrices, which ensures perfect recovery in the random case, i.e., that $(U_{1},V_{1})$ sufficiently more dense than all other $m_{1}\times n_{1}$ subgraphs.
@@ 2.3.2 Perfect Recovery Under the Adversarial Planted Submatrix Model
  • adding $\Omega(m_{1})$ entries within $\boldsymbol{A}[U_{1},V-V_{1}]$ and/or $\Omega(n_{1})$ entries within $\boldsymbol{A}[U-U_{1},V_{1}]$.
@@ 2.3.3 Relationship with Existing Recovery Guarantees
The sufficient conditions for recovery given by Theorem 2.1 and Theorem 2.2 can be shown to be extensions of several existing sufficient conditions for perfect recovery of dense subgraphs, planted cliques, and dense submatrices. Indeed, consider the $K=2$-case and set
@@
In this case, the sufficient condition given in Theorem 2.1 immediately specializes to that given in (Bombina and Ames, 2020, Theorem 3.1). Moreover, placing further assumptions on $q$ and $p$, e.g., $q=1$, yield the recovery guarantees for the planted clique and maximum edge biclique problem proposed in Ames and Vavasis (2011). Simiilarly, the sufficient conditions for perfect recovery under the adversarially setting given in Theorem 2.2 directly specialize the sufficient conditions for perfect recovery of planted cliques and bicliques given in (Ames and Vavasis, 2011, Section 4.1 and Section 5.1).

### v2

@@ Abstract
We consider the densest submatrix problem, which seeks the submatrix of fixed size of a given binary matrix that contains the most nonzero entries. This problem is a natural generalization of fundamental problems in combinatorial optimization, e.g., the densest subgraph, maximum clique, and maximum edge biclique problems, and has wide application in the study of complex networks. Much recent research has focused on the development of sufficient conditions for exact solution of the densest submatrix problem via convex relaxation. The vast majority of these sufficient conditions establish identification of the densest submatrix within a graph containing exactly one large dense submatrix hidden by noise. The assumptions of these underlying models are not observed in real-world networks, where the data may correspond to a matrix containing many dense submatrices of varying sizes.
@@ 1 Introduction
We consider the densest submatrix problem (DSM), in which one aims to identify the submatrix of given size containing a maximum number of nonzero entries; see Lanciano et al. (2024) for a survey of the historical and recent study of the DSM. The densest submatrix can be considered a generalization of the densest bipartite subgraph problem: given a bipartite graph $G=(V,U,E)$ and integers $k_{1}$ and $k_{2}$, the objective is to extract a subgraph induced by $k_{1}$ nodes from $V$ and $k_{2}$ nodes from $U$ that maximizes the number of edges. This further generalizes the densest subgraph problem: given graph $G$ and integer $k$, find the $k$-node subgraph containing maximum number of edges.
@@ 1.1 Contributions
In this work, the phrase “with high probability” (abbreviated as w.h.p.) refers to events whose probability converges to 1 as the problem size increases. Formally, for a sequence of random events $\{E_{n}\}$, we say that $E_{n}$ occurs with high probability if $\Pr(E_{n})\to 1$ polynomially as $n\to\infty$ in the sense that $\Pr(E_{n})\geq 1-1/p(n)$ for some polynomial function $p(n)$ with $p(n)\to\infty$ as $n\to\infty$.
@@
for each $i\in U_{r}$, $j\in V_{s}$, $r\in\{1,2,\dots,k_{1}\}$ and $s\in\{1,2,\dots,k_{2}\}$ for probability matrix $\boldsymbol{P}=[p_{rs}]\in\mathbb{R}^{k_{1}\times k_{2}}$ with values in the interval $[0,1]$.
@@
$\tilde{\sigma}^{2}:=\frac{p^{*}}{1-p^{*}}\text{ where }p^{*}:=\max_{(r,s)\neq(1,1)}p_{rs}.$
@@
$p_{11}-p^{*}$ $\geq c\cdot\max\left\{\sqrt{\frac{\tilde{\sigma}^{2}N\log N}{m^{2}}},\sqrt{\frac{\sigma_{11}^{2}\log N}{m}},\sqrt{\frac{\max\{\tilde{\sigma}^{2},1\}(\log N)^{3}}{m}}\right\}.$ (1b)
@@
Here, we denote $(U_{1},V_{1})$ as the indices of the block of interest for notational simplicity and consistency through the rest of the manuscript. We assume that the planted block is indexed by $U_{1},V_{1}$ without loss of generality, i.e., the result holds if $(U_{i},V_{j})$ indexes a planted block satisfying (1b) (with $p^{*}$ redefined as the maximum probability over all blocks except the $(i,j)$ block), for any $i,j\in\{1,2,\dots,k\}$. In this case, $(U_{i},V_{j})$ indexes the densest submatrix of $\boldsymbol{A}$ with high probability.
This result generalizes to the unbalanced planted submatrix model, as well as adversarially constructed matrices. Note that Theorem 1.1 omits an explicit description of the family of convex relaxations that yield the densest submatrix w.h.p. under the hypothesis of Theorem 1.1. We will derive the relevant convex relaxation and provide these generalized sufficient conditions in the following section, including a range of regularization parameters for which the relaxation is exact. Specifically, we provide a generalization of Theorem 1.1 in Theorem 2.1. The corresponding result for the adversarial case is found in Theorem 2.2.
@@ 2.1 Formulation and Relaxation
We define $M]:=\{1,2,...,M\}$ and $N]:=\{1,2,...,N\}$ for every positive integer $M$ and $N$ respectively. Let $\bar{U}_{1}\subseteq[M]$ and $\bar{V}_{1}\subseteq[N]$ denote the row and column indices of a submatrix of $\boldsymbol{A}$ such that $|\bar{U}_{1}|=m$ and $|\bar{V}_{1}|=n$. Furthermore, we define vectors $\bar{\boldsymbol{u}}\in\{0,1\}^{M}$ and $\bar{\boldsymbol{v}}\in\{0,1\}^{N}$ as the indicator vectors of $\bar{U}_{1}$ and $\bar{V}_{1}$ respectively. The support of this submatrix is given by the nonzero entries of the matrix $\bar{\boldsymbol{X}}=\bar{\boldsymbol{u}}\bar{\boldsymbol{v}}^{\top}\in\{0,1\}^{M\times N}$. By construction, $\bar{\boldsymbol{X}}=\bar{\boldsymbol{u}}\bar{\boldsymbol{v}}^{\top}$ is a rank-one binary matrix supported on $\bar{U}_{1}\times\bar{V}_{1}$, such that $\bar{X}_{ij}=1$ if $(i,j)\in\bar{U}_{1}\times\bar{V}_{1}$ and $\bar{X}_{ij}=0$ otherwise.
Suppose that $\Omega$ encodes the set of zero entries in a given matrix $\boldsymbol{A}\in\mathbb{R}^{M\times N}$. Without loss of generality, we may assume that $\boldsymbol{A}$’s entries are binary. If not, we can substitute $\boldsymbol{A}$ with a binary matrix that features the same sparsity pattern, preserving the index set of the densest $m\times n$-submatrix. Our objective is to find a rank-one matrix $\boldsymbol{X}$ with $mn$ nonzero entries that minimizes the number of disagreements with $\boldsymbol{A}$ on $\Omega$.
@@
$d(\bar{\boldsymbol{X}})=\frac{1}{mn}\left(mn-\|P_{\Omega}(\bar{\boldsymbol{X}})\|_{0}\right),$
@@
$\begin{array}[]{cl}\displaystyle\min_{\boldsymbol{X},\boldsymbol{Y}\in\{0,1\}^{M\times N}}&\mathrm{Tr}(\boldsymbol{Y}\boldsymbol{1}\boldsymbol{1}^{T})\\ \mathrm{s.t.}&\mathrm{Tr}(\boldsymbol{X}\boldsymbol{1}\boldsymbol{1}^{T})=mn,~P_{\Omega}(\boldsymbol{X}-\boldsymbol{Y})=\boldsymbol{0},~\mathrm{rank}(\boldsymbol{X})=1.\end{array}$ (2)
@@
We move the rank constraint to the objective and relax using the nuclear norm Recht et al. (2010), and relax binary constraints using appropriate box constraints. This yields the optimization problem
$\begin{array}[]{cl}\min&\lVert\boldsymbol{X}\rVert_{*}+\gamma\mathrm{Tr}(\boldsymbol{Y}\boldsymbol{1}\boldsymbol{1}^{T})\\ \mathrm{s.t.}&\mathrm{Tr}(\boldsymbol{X}\boldsymbol{1}\boldsymbol{1}^{T})=mn,~P_{\Omega}(\boldsymbol{X}-\boldsymbol{Y})=\boldsymbol{0},~\boldsymbol{0}\leq\boldsymbol{X}\leq\boldsymbol{1}\boldsymbol{1}^{T},~\boldsymbol{0}\leq\boldsymbol{Y}.\end{array}$ (3)
Here, $\gamma\geq 0$ denotes a regularization parameter selected to balance between the two objectives. and $\|\boldsymbol{X}\|_{*}$ denotes the nuclear norm of matrix $\boldsymbol{X}$, which is equal to the sum of the singular values of $\boldsymbol{X}$. The relaxation (3) is convex since it is the minimization of the sum of a norm and linear function subject to linear constraints. Moreover, (3) satisfies Slater’s condition; indeed, $\boldsymbol{X}=\boldsymbol{Y}=m_{1}n_{1}/(MN)\boldsymbol{1}\boldsymbol{1}^{T}$ is strictly feasible for  (3). Moreover, (3) can be rewritten as a semidefinite program using the argument presented in Recht et al. (2010). Thus, (3) can be solved using standard algorithms for semidefinite programming such as interior point methods; we consider a first-order method for solution of (3) in Section 5.
As previously mentioned, it is reasonable to anticipate retrieving the solution of (2) from (3) when matrix $\boldsymbol{A}$ consists of a dense $m_{1}\times n_{1}$ submatrix hidden among perhaps many other dense submatrices under certain assumptions about the size and density of these submatrices. In particular, this program succeeds in recovering the planted block exactly when embedded in a noisy binary matrix under the conditions of Theorem 1.1.
@@ 2.2 A Sufficient Condition for Perfect Recovery for Random Matrices
Let $\boldsymbol{A}\in\mathbb{R}^{M\times N}$ be sampled from the planted submatrix model with row partitions $(U_{1},\ldots,U_{k})$ and column partitions $(V_{1},\ldots,V_{k})$. Let $m_{i}=|U_{i}|$ and $n_{i}=|V_{i}|$, and let $p_{rs}$ denote the probability that $A_{ij}=1$ for $i\in U_{r}$, $j\in V_{s}$.
@@
$\tilde{\sigma}^{2}:=\max_{(r,s)\neq(1,1)}\left\{\frac{p_{rs}}{1-p_{rs}}\right\}.$
Then there exist constants $c_{1},c_{2}>0$ such that the planted block $(U_{1},V_{1})$ is the unique densest $m_{1}\times n_{1}$ submatrix of $\boldsymbol{A}$ and $\boldsymbol{X}^{*}=\boldsymbol{u}_{1}\boldsymbol{v}_{1}^{T}$ unique optimal solution of the convex relaxation (3) with high probability if the following conditions hold for $(r,s)\neq(1,1)$
@@
$p_{11}-p^{*}$ $\geq\max\left\{\sqrt{\frac{\tilde{\sigma}^{2}N^{*}\log N^{*}}{m_{1}n_{1}}},\sqrt{\frac{\sigma_{11}^{2}\log N^{*}}{\min\{m_{1},n_{1}\}}},\sqrt{\frac{\max\{\tilde{\sigma}^{2},1\}(\log N^{*})^{3}}{\min\{m_{1},n_{1}\}}}\right\}$ (4c)
@@
$\gamma\in\left(\frac{2}{(p_{11}-p^{*})\sqrt{m_{1}n_{1}}},\frac{c_{2}}{(p_{11}-p^{*})\sqrt{m_{1}n_{1}}}\right),$ (5)
@@
Remark 2.
The assumption that $m_{1}n_{1}\leq m_{r}n_{s}$ for all $(r,s)\neq(1,1)$ is made to simplify both the statement of the sufficient conditions for perfect recovery and their proof. It is reasonable to expect perfect recovery when $(U_{1},V_{1})$ is sufficiently more dense in expectation than any other $m_{1}\times n_{1}$ block of $\boldsymbol{A}$ and $m_{1}\times n_{1}$ is sufficiently large. Assumptions (4a) and (4c) characterize when this is satisfied in Theorem 2.1. An analogous result holds when we replace Assumption (4a) with the assumption that $p_{11}$ exceeds the average expected density of any union of blocks indexing a submatrix with at least $m_{1}n_{1}$ entries of $\boldsymbol{A}$, albeit with far more complicated formal statement and notation than that given in Theorem 2.1. We give a deterministic characterization of this more general sufficient condition in the next section.
@@ 2.3 A Sufficient Condition for Perfect Recovery for Adversarially Generated Matrices
The sufficient conditions for perfect recovery given in Section 2.2 may be overly conservative. As note in, the assumption that the planted submatrix of interest is smaller than all other blocks in the generative model is kept largely to ensure that this is the densest $m_{1}\times n_{1}$-submatrix with high probability if the hypothesis of Theorem 2.1 is satisfied. This simplifies the statement of the sufficient conditions, as well as their proof, but is not necessary, in general, to ensure perfect recovery. In this section, we provide a sufficient condition for recovery that holds for deterministically generated graphs. This condition deterministically encodes the relationship between planted submatrices, which ensures perfect recovery in the random case, i.e., that $(U_{1},V_{1})$ sufficiently more dense than all other $m_{1}\times n_{1}$ subgraphs.
@@ 2.3.2 Perfect Recovery Under the Adversarial Planted Submatrix Model
  • adding $\Omega(m_{1})$ entries within a column of $\boldsymbol{A}[U_{1},V-V_{1}]$ and/or $\Omega(n_{1})$ entries within a row of $\boldsymbol{A}[U-U_{1},V_{1}]$.
@@ 2.3.3 Relationship with Existing Recovery Guarantees
The sufficient conditions for recovery given by Theorem 2.1 and Theorem 2.2 can be shown to be extensions of several existing sufficient conditions for perfect recovery of dense subgraphs, planted cliques, and dense submatrices. Indeed, consider the case when $K=2$ and set
@@
In this case, the sufficient condition given in Theorem 2.1 immediately specializes to that given in (Bombina and Ames, 2020, Theorem 3.1). Moreover, placing further assumptions on $q$ and $p$, e.g., $q=1$, yield the recovery guarantees for the planted clique and maximum edge biclique problem proposed in Ames and Vavasis (2011). Similarly, the sufficient conditions for perfect recovery under the adversarially constructed setting given in Theorem 2.2 directly specialize the sufficient conditions for perfect recovery of planted cliques and bicliques given in (Ames and Vavasis, 2011, Section 4.1 and Section 5.1).
Our results are closely related to similar recovery guarantees for biclustering under generalizations of the stochastic block model. Specifically, several recent results Ames (2014); Sudoso (2025) establish sufficient conditions under which we can find all planted dense submatrices, corresponding to biclusters in the bipartite graph with adjacency matrix $\boldsymbol{A}$, for $\boldsymbol{A}$ sampled from the planted submatrix model; the planted submatrix model corresponds to a special case of the stochastic block model (SBM) here. In this context, Theorem 2.1 and Theorem 2.2 give sufficient conditions for perfect recovery of a single planted dense submatrix for $\boldsymbol{A}$ sampled from the planted submatrix model. This translates to a sufficient condition for perfect recovery of the single densest bicluster of a given size in the biclustering setting. Comparing these results suggests that there are instances sampled from the planted submatrix model where the planted dense $m\times n$-submatrix corresponds to one of the biclusters in the corresponding SBM. In this case, biclustering should reveal the densest $m\times n$-submatrix, along with all other biclusters. In this sense, we can view the densest submatrix problem as that of identifying a single hidden dense bicluster or submatrix, among potentially many biclusters with heterogeneous edge probabilities and outliers.
The sufficient condition for solution of the $m\times n$-densest submatrix problem includes a broader range of matrices than the sufficient conditions for perfect recovery of biclusters given in Ames (2014); Sudoso (2025) in the following ways:
  • The sufficient conditions given in Theorem 2.1 hold under a larger number of outliers than those for biclusters. Indeed, sufficient conditions for perfect biclustering under the SBM restrict the number of outlier nodes to $\mathcal{O}(\min\{\min\{m_{i},n_{i}\}\})$, where $(m_{i},n_{i})$ is the size of the $i$th planted bicluster. The sufficient conditions given in Theorem 2.1 ensure recovery of the densest $m\times n$-submatrix for instances $\boldsymbol{A}\in\{0,1\}^{M\times N}$ of the planted submatrix model with up to $\mathcal{O}(\min\{M,N\})$ outliers nodes.
  • Moreover, our construction of the planted submatrix model potentially allows multiple relatively dense blocks in each row or column of blocks, while the stochastic block model assumes one especially dense block per row and column (typically indexed as the diagonal block).
  • Solution of (3) requires selection of the submatrix sizes $(m,n)$, while most biclustering methods do not need to be provided block sizes. However, many biclustering methods, require the number of biclusters $k$ as a parameter, which is not needed by (3).
This implies that we are able to recover the densest $m\times n$-submatrix using the convex relaxation (3) in a wider range of instances $\boldsymbol{A}$ than if we identify all biclusters, including the densest $m\times n$-submatrix, in the bipartite graph with adjacency matrix $\boldsymbol{A}$.